For each of the following cases, calculate the point price elasticity of demand, and state whether demand is elastic, inelastic, or unit elastic. The demand curve is given by

QD = 5,000 – 50 PX

a. If the price of the product is $50.

b. If the price of the product is $75.

2) For each case, should the firm raise or lower the prices to maximize revenues? Why or Why not? Explain.

3) Suppose the income elasticity of demand is 1.5.

What impact will a recession cause for the demand for your product? Why?   Explain.

4) The cross-price elasticity of demand between your good and a related good is -3.5.

What can you determine about consumer demand for your product from this information?   Suppose the price of the related good changes by 10%?   How will this affect your sales?

Explain.
Answer #1

Cunul is given by! Demand QD - 5020 - 50 PX Now point elasticity of demand Ed = a) Now, QD = 5000-50 px dos y en los - dab =-tile ed = -50* ( 75 ) ) Now, absolute value of As led l = 3 which is elasticity led 1=3 greater than 1, hence demand is elast2 As in come elasticity of demand is ! which means that income elasticity of a more than 1. Hace demand is in come. Which mea